# justification=centering

### v1

@@ Appendix B Pixel correlation between datasets
To study the correlation between datasets, we adopted the approach used in 2024A&A...688A.185J, using a correlation metric from the VIP package based on the Spearman rank correlation coefficient. This coefficient was chosen to avoid the normality assumption required by the Pearson coefficient and to circumvent the dependence of the Structural Similarity Index (SSIM) on the image dynamic range, which will be normalized later in the post-processing. For each dataset, we identified the frame within the science cube that exhibits the highest correlation with other frames, designating it as the representative frame of that dataset. We then computed the pixel correlation between this representative frame and the representative frames of all other datasets in the cluster. This process was repeated for all datasets within one cluster as well as for all clusters. The dataset exhibiting the highest correlation with the other datasets in its cluster is expected to share the most prominent noise features and thus be the most representative. We then compared the Spearman correlation of the selected cluster centers, chosen based on environmental conditions, to that of the most correlated dataset in each cluster. The results are shown in Figure 20.
The high correlation values for the cluster centers confirm that these selections reliably represent their respective clusters. This correlation metric proves less effective for clusters affected by strong wind-driven halo effects (clusters 4 and 5). This limitation arises from the rotation of wind-driven speckles with the parallactic angle (2020A&A...638A..98C), which introduces additional complexity. In such cases, high correlation coefficients tend to identify datasets with less pronounced wind effects rather than true feature similarity, reducing the effectiveness of correlation as a metric in these specific scenarios.
@@
Appendix C Clustering illustration using principal components
To analyze the distinctions between clusters, we projected the environmental datasets onto their first three principal components. Figure 21 (top) shows that most clusters are well-separated, although clusters 2 and 6 overlap. This overlap is resolved when examining the third principal component, as shown in Plot 21 (bottom).
Figure 21: Visualization of clusters using principal component analysis. Top shows the first vs second component projection. Bottom shows second vs third component projection.
Appendix D Optimal parameters for cluster centers
Table 2: Optimal parameters for PSF subtraction techniques and RSM for each cluster center. The RSM parameters include the planetary flux multiplicative factor ($\delta$), defined as a multiple of the standard deviation, the crop size, and the variance term used to determine the noise estimation region within the annulus. The noise estimation methods are denoted as follows: ”SM” (Segment with mask-based estimation), ”ST” (Spatio-Temporal estimation), and ”FR” (Frame-based estimation) (for further details, see 2021A&A...656A..54D).
@@ Appendix D Optimal parameters for cluster centers
Appendix E Variation of optimal RSM parameters across cluster centers and individual datasets
@@
Appendix F Comparison with the Dahlqvist contrast curve definition
This section compares the contrast curve methodology used in 2022A&A... [^6] A..33D for the SHARDDS survey with the approach applied in our study. 2022A&A... [^6] A..33D define the detection threshold based on the first false positive within the field of view and considers 95% recovery of injected signals. In contrast, our approach employs a detection threshold corresponding to an FAP of $3\times 10^{-7}$ under a lognormal distribution at a given separation and evaluates contrast at the 50% injection recovery level. Figure 23 presents these contrast curves for two datasets: one without a potential signal (Fig. 23(a)) and another with a potential signal (Fig. 23(b)). In Fig. 23(a), the contrast curves derived from different detection thresholds remain consistent at the same completeness levels. However, this trend does not hold in Fig. 23(b), where the contrast curve based on the lognormal detection threshold outperforms that derived from the first false positive at both completeness levels. In both figures, there is only a minor gap between the contrast levels required for 50% and 95% signal recovery. This result highlights a key limitation of the approach in 2022A&A... [^6] A..33D, where the contrast is influenced by the brightest signal in the field of view without determining whether it originates from noise or a potential candidate —a limitation that also applies to thresholds based on the maximum F1 score. Moreover, this method defines the detection threshold based on the entire field of view rather than at each separation, limiting its ability to assess contrast variations across different separations.
Figure 23: Comparison of contrast curves for two datasets, one with no potential companion (a) and one with a potential companion (b): 95% completeness using the first false positive threshold and lognormal threshold (purple and green), and 50% completeness for both thresholds (red and blue)
Appendix G Comparing RSM noise distribution models
@@
Appendix H Maximum F1 score threshold generalized across a cluster
@@
This section evaluates the generalization of thresholds determined using the maximum F1 score from a given dataset across other datasets with similar observing conditions. The maximum F1 score is highly sensitive to false positives, often influenced by the brightest speckle in the representative dataset rather than the overall noise profile. Generalizing such thresholds can lead to inconsistencies, as the presence of these speckles reflects specific environmental parameters and their interaction with adaptive optics. These factors are inherently unique to individual observations and cannot be reliably generalized, even under broadly similar observing conditions. Figure 25 illustrates this issue by generalizing the maximum F1 score threshold from the cluster center BD-15 705 to the dataset HIP 92680 in cluster 1. The threshold derived from the cluster center (pink) shows a factor of 10 difference compared to the self-computed maximum F1 score ( cyan) and the lognormal threshold (dark blue). This significant discrepancy leads to the generalized threshold accepting false positives that would otherwise be excluded, underscoring the difficulties in applying overly tuned thresholds across datasets.
Appendix I Comparison between RSM maps and PCA Signal-to-Noise Ratio maps
@@
Figure 26: Comparison of RSM and PCA S/N maps for several newly identified background stars. The left panels show the RSM maps generated with the APCA–NMF–LOCI–FMKLIP combination in the H2 and H3 filters, along with their corresponding RSM scores. The right panels display the PCA S/N maps computed with 5 components.

### v2

@@ Appendix A Pixel correlation between datasets
To study the correlation between datasets, we adopted the approach used in Juillard et al. (2024), using a correlation metric from the VIP package based on the Spearman rank correlation coefficient. This coefficient was chosen to avoid the normality assumption required by the Pearson coefficient and to circumvent the dependence of the Structural Similarity Index (SSIM) on the image dynamic range, which will be normalized later in the post-processing. For each dataset, we identified the frame within the science cube that exhibits the highest correlation with other frames, designating it as the representative frame of that dataset. We then computed the pixel correlation between this representative frame and the representative frames of all other datasets in the cluster. This process was repeated for all datasets within one cluster as well as for all clusters. The dataset exhibiting the highest correlation with the other datasets in its cluster is expected to share the most prominent noise features and thus be the most representative. We then compared the Spearman correlation of the selected cluster centers, chosen based on environmental conditions, to that of the most correlated dataset in each cluster. The results are shown in Figure 20.
The high correlation values for the cluster centers confirm that these selections reliably represent their respective clusters. This correlation metric proves less effective for clusters affected by strong wind-driven halo effects (clusters 4 and 5). This limitation arises from the rotation of wind-driven speckles with the parallactic angle (Cantalloube et al. 2020a), which introduces additional complexity. In such cases, high correlation coefficients tend to identify datasets with less pronounced wind effects rather than true feature similarity, reducing the effectiveness of correlation as a metric in these specific scenarios.
@@
Appendix B Clustering illustration using principal components
To analyze the distinctions between clusters, we projected the environmental datasets onto their first three principal components. Figure 21 (top) shows that most clusters are well-separated, although clusters two and 6 overlap. This overlap is resolved when examining the third principal component, as shown in Plot 21 (bottom).
Figure 21: Visualization of clusters using principal component analysis. Top: First versus second component projection. Bottom: Second versus third component projection.
Appendix C Optimal parameters for cluster centers
Table 1: Optimal parameters for PSF subtraction techniques and RSM for each cluster center. [^4] The RSM parameters include the planetary flux multiplicative factor ($\delta$), defined as a multiple of the standard deviation, the crop size, and the variance term used to determine the noise estimation region within the annulus. The noise estimation methods are denoted as follows: ”SM” (Segment with mask-based estimation), ”ST” (Spatio-Temporal estimation), and ”FR” (Frame-based estimation) (for further details, see Dahlqvist et al. 2021a).
@@ Appendix C Optimal parameters for cluster centers
Appendix D Variation of optimal RSM parameters across cluster centers and individual datasets
@@
Appendix E Comparison with the Dahlqvist contrast curve definition
This section compares the contrast curve methodology used in Dahlqvist et al. (2022) for the SHARDDS survey with the approach applied in our study. Dahlqvist et al. (2022) define the detection threshold based on the first false positive within the field of view and considers 95% recovery of injected signals. In contrast, our approach employs a detection threshold corresponding to an FAP of $3\times 10^{-7}$ under a lognormal distribution at a given separation and evaluates contrast at the 50% injection recovery level. Figure 23 presents these contrast curves for two datasets: one without a potential signal (Fig. 23(a)) and another with a potential signal (Fig. 23(b)). In Fig. 23(a), the contrast curves derived from different detection thresholds remain consistent at the same completeness levels. However, this trend does not hold in Fig. 23(b), where the contrast curve based on the lognormal detection threshold outperforms that derived from the first false positive at both completeness levels. In both figures, there is only a minor gap between the contrast levels required for 50% and 95% signal recovery. This result highlights a key limitation of the approach in Dahlqvist et al. (2022), where the contrast is influenced by the brightest signal in the field of view without determining whether it originates from noise or a potential candidate —a limitation that also applies to thresholds based on the maximum F1 score. Moreover, this method defines the detection threshold based on the entire field of view rather than at each separation, limiting its ability to assess contrast variations across different separations.
Figure 23: Comparison of contrast curves for two datasets, one with no potential companion (a) and one with a potential companion (b). Contrasts are generated with 95% completeness using the first false positive threshold and lognormal threshold (in purple and green), and 50% completeness for both thresholds (in red and blue)
Appendix F Comparing RSM noise distribution models
@@
Appendix G Maximum F1 score threshold generalized across a cluster
@@
This section evaluates the generalization of thresholds determined using the maximum F1 score from a given dataset across other datasets with similar observing conditions. The maximum F1 score is highly sensitive to false positives, often influenced by the brightest speckle in the representative dataset rather than the overall noise profile. Generalizing such thresholds can lead to inconsistencies, as the presence of these speckles reflects specific environmental parameters and their interaction with adaptive optics. These factors are inherently unique to individual observations and cannot be reliably generalized, even under broadly similar observing conditions. Figure 25 illustrates this issue by generalizing the maximum F1 score threshold from the cluster center BD-15 705 to the dataset HIP 92680 in cluster 1. The threshold derived from the cluster center (pink) shows a factor of ten difference compared to the self-computed maximum F1 score ( cyan) and the lognormal threshold (dark blue). This significant discrepancy leads to the generalized threshold accepting false positives that would otherwise be excluded, underscoring the difficulties in applying overly tuned thresholds across datasets.
Appendix H Comparison between RSM maps and PCA signal-to-noise ratio maps
@@
Figure 26: Comparison of RSM and PCA S/N maps for several newly identified background stars. Left and middle column panels: RSM maps generated with the APCA–NMF–LOCI–FMKLIP combination in the H2 and H3 filters, along with their corresponding RSM scores. Right column panels: PCA S/N maps computed with five components.